# Complexity reduction of physical models: an equation-free approach by means of scaling

### v1

@@ 5 A Case Study: the classical projectile model
First, we use the Theorems from Appendix B to formulate the projectile model with the minimal number of parameters. Properties A.4, A.5 and A.6 are satisfied by the $N_{\mathsf{p}}=4$ physical parameters $\tilde{p}_{1},\tilde{p}_{2},\tilde{p}_{3},\tilde{p}_{4}$ (6), with
$[\tilde{p}_{1}]=[\tilde{g}]=\mathrm{m}\,\mathrm{s}^{-2},\quad[\tilde{p}_{2}]=[\tilde{R}]=\mathrm{m},\quad[\tilde{p}_{3}]=[\tilde{h}_{0}]=\mathrm{m},\quad[\tilde{p}_{4}]=[\tilde{v}_{0}]=\mathrm{m}\,\mathrm{s}^{-1},$ (23)
being $\mathrm{m}$ meters and $\mathrm{s}$ seconds. Then, Theorem B.3 guarantees that there exist only $N_{\mathsf{0}}=N_{\mathsf{p}}-\mathrm{rank}(\textbf{M})=2$ independent dimensionless parameters, where M is the matrix
@@
and the $N_{\mathsf{0}}=2$ independent dimensionless parameters can be computed as in (B.12), i.e.,
@@
Properties A.1 and A.2 are fulfilled by $N_{\mathsf{l}}=4$ coefficients $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ (5), with
@@
Then, by also assuming Property A.3 for $\tilde{\theta}_{1},\tilde{\theta}_{2}$ (3), i.e.,
@@
Theorem B.4 guarantees that the coefficients $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ (5) can be written as power-law monomials of $\pi_{1},\pi_{2}$ (25) only, i.e.,
@@
Finally, Theorem B.6 indicates under Properties A.1, A.2, A.3, A.4, A.5 and A.6 that the vector
@@
of exponents in (29) must satisfy the linear system $\textbf{S}\,\vec{\alpha}=\vec{s}$ (B.40), which, for the projectile model, reads in row echelon form as
@@ 5.1 Optimally Scaled Coefficients
Theorem B.7 allows solving the optimization problem (20) for any fixed choice of $\pi_{1},\pi_{2}$ (25), since one can set (31) as the constraint $\boldsymbol{\Sigma}\,\vec{\alpha}=\vec{\sigma}$ (B.48) to satisfy the assumption (B.50). In particular, the system (B.49) can be solved to find optimal values of $\vec{\alpha}$ (30) for plugging into (29). Then, the system (B.49) reads as
@@ 5.2 Generalized Optimal Scaling $(\mathtt{s}=0)$: Results
Table 1 states that such coefficients allow minimizing the variability $\mathop{\mathrm{max}}_{i}\lambda_{i}/\mathop{\mathrm{min}}_{i}\lambda_{i}$. In order to verify such a claim, we compare the values achieved by coefficients (35) with the obtained ones by the following approach suggested in the literature [13, 14]. Provided the values of physical parameters $\tilde{p}_{1},\tilde{p}_{2},\tilde{p}_{3},\tilde{p}_{4}$ (6), it is possible to impose $N_{\mathsf{x}}=2$ out of the $N_{\mathsf{l}}=4$ coefficients $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ (5) to be equal to $1$. Then, one aims to solve the resulting system of equations with respect to $\tilde{\theta}_{1}$ and $\tilde{\theta}_{2}$. Finally, the corresponding values of $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ (5) are obtained by plugging the computed $\tilde{\theta}_{1}$ and $\tilde{\theta}_{2}$ into (5).
@@
The Matlab solver $\mathrm{\it{ode45}}$ [16] is employed to compute the numerical solution $\varphi=\varphi_{n}$ to the Ordinary Differential Equation (4) in the time interval $[0,T/\theta_{1}]$. The subscript $n$ indicates that the numerical values of coefficients $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ and any computed quantity at each time step are rounded to $n$ decimal places. The relative error is estimated as
[…]
where $\varphi_{\mathrm{ref}}$ is the numerical solution to the Ordinary Differential Equation (4) with $\lambda_{1}=g$, $\lambda_{2}=R^{-1}$, $\lambda_{3}=h_{0}$, $\lambda_{4}=v_{0}$, i.e., $\theta_{1}=\theta_{2}=1$, that has been taken as a reference in the time interval $[0,T]$. Such a solution $\varphi_{\mathrm{ref}}$ is obtained by using the Matlab solver $\mathrm{\it{ode45}}$ [16] with the default 16 significant digits of precision for numerical computations. A relative error tolerance of $10^{-5}$ and an absolute error tolerance of $10^{-8}$ have been required for all the computations performed by $\mathrm{\it{ode45}}$ [16], being $2$ orders of magnitude smaller than the default tolerances. The Simpson’s quadrature rule [17] has been used to compute the integral in (40).
@@ 5.3 Generalized Optimal Scaling $(\mathtt{s}=1,2,3,4)$: Results
$T_{0}$$\coloneqq\begin{cases}\mathop{\mathrm{min}}\left\{\tau\geq 0\,:\,\left|1\,+\,\lambda_{2}\,\varphi_{\mathrm{ref}}(\tau)\right|\leq 10^{-2}\right\},&\mbox{if}\quad\mathtt{s}=1,2,\\ \mathop{\mathrm{min}}\left\{\tau\geq 0\,:\,\left|1\,+\,\lambda_{2}\,\varphi_{\mathtt{s}}(\tau)\right|\leq 10^{-2}\,\lor\,\left|1\,+\,\lambda_{2}\,\varphi_{\mathrm{ref}}(\tau)\right|\leq 10^{-2}\right\},&\mbox{if}\quad\mathtt{s}=3,4,\\ \end{cases}$ (42)
where $\varphi_{\mathtt{s}}(\tau)$ is the approximating solution (41), while $\varphi_{\mathrm{ref}}(\tau)$ is the numerical solution to (4) provided by the Matlab solver $\mathrm{\it{ode45}}$ [16], that has been taken as a reference since it is the solution one aims to approximate. If $\mathtt{s}=3,4$, the approximating solution (41) is also numerically computed by means of the Matlab solver $\mathrm{\it{ode45}}$ [16]. A relative error tolerance of $10^{-6}$ and an absolute error tolerance of $10^{-9}$ have been required for all the computations performed by $\mathrm{\it{ode45}}$ [16], i.e., $3$ orders of magnitude smaller than the default tolerances. The Simpson’s quadrature rule [17] has been used to compute the integrals in (42). The time $T$ in (42) is chosen with the aim of stopping the numerical integration just before singularities are attained in (4) or (41). Such irregularities appear when the computed solutions of (4) or (41) are negative, i.e., equal to $-1/\lambda_{2}<0$. In other words, the time interval $[0,T]$ in (42) allows investigating the dynamics of the projectile further than it reaches the ground level. The resulting error $\epsilon_{\mathtt{s}}(T)$ (42) is shown in Figure 4 as a function of $\pi_{1}$ and $\pi_{2}$ (25).
@@ 6 Conclusions & Discussion
We reviewed methodologies for scaling of dimensional models and presented a novel approach which we called Generalized Optimal Scaling, or GOS. GOS inherits all the features of its predecessors, i.e., Optimal Scaling (OS) and Optimal Scaling with Constraints (OSC), including the ability to quantify conditions for asymptotic models and, in addition, provides control of over-parametrization - the property beneficial for complex multiscale models. Both OS and OSC can be recovered from GOS with the special choices of the GOS’ parameter $\mathtt{s}$. Such a non-negative integer $\mathtt{s}$ is the index of the coefficient $\lambda_{\mathtt{s}}$ vanishing to 00 when an asymptotic model is considered. The GOS method was rigorously formulated and carefully tested on the classical projectile model.
By setting $\mathtt{s}=0$ in GOS (see Section 4), one makes use of the Optimal Scaling methodology [6] presented in Section 3.1. For the projectile model, we demonstrated that OS allows reducing the variability of the involved coefficients by at least two more orders of magnitude compared to other scaling approaches traditionally applied to this model. Such a reduction is beneficial for diminishing round-off and numerical errors. Indeed, we showed that the coefficients provided by the Optimal Scaling approach allow reducing round-off errors by two more orders of magnitude compared to the traditional scaling techniques.
When GOS is applied with $\mathtt{s}\geq 1$ (see Section 4), one performs the Optimal Scaling with Constraints [7] explained in Section 3.2. Such an approach provides a threshold on values of parameters that allow discarding terms in the considered equations, without significantly modifying the solutions. Thus, one can drop some terms and consider a simpler model, if the provided criterion is met. For the projectile model, OSC is able to identify values of parameters yielding to a relative approximation error of at most $10\%$ for all simulated intervals of time and all $\mathtt{s}\geq 1$.
A natural future step is an incorporation of GOS in an automated tool to streamline model reduction and scaling, optimizing parameter selection before further modeling steps. This would facilitate the use of GOS in complex, high-dimensional models, even for users with less mathematical expertise.
@@ Acknowledgements
Appendix A Properties
The Optimal Scaling approaches presented in this paper rely on the Properties listed below. Such assumptions give rise to the Theorems formulated in Appendix B to constitute the tools for Optimal Scaling procedures.
Property A.1.
$\kappa_{i},\theta_{j}\in(0,\infty),\quad c_{i,j}\in\mathbb{R},\quad\forall i=1,\dots,N_{\mathsf{l}},\quad\forall j=1,\dots,N_{\mathsf{x}},$ (A.1)
where each $\tilde{\kappa}_{i}$ is an explicit function of $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$ only, and all the exponents $c_{i,j}$ in (A.1) are explicitly known real numbers.
Property A.2.
Each of the factors $\tilde{\kappa}_{1},\dots,\tilde{\kappa}_{N_{\mathsf{l}}}$ in (A.1) can be written as a power-law monomial of the physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$ only:
$\tilde{\kappa}_{i}=\tilde{p}_{1}^{d_{i,1}}\,\tilde{p}_{2}^{d_{i,2}}\,\cdots\,\tilde{p}_{N_{\mathsf{p}}}^{d_{i,N_{\mathsf{p}}}},\quad d_{i,j}\in\mathbb{R},\quad\forall i=1,\dots,N_{\mathsf{l}},\quad\forall j=1,\dots,N_{\mathsf{p}},$ (A.2)
with all the exponents $d_{i,j}$ in (A.2) to be explicitly known real numbers.
Property A.3.
$\tilde{\theta}_{j}=\tilde{p}_{1}^{t_{1,j}}\,\tilde{p}_{2}^{t_{2,j}}\,\cdots\,\tilde{p}_{N_{\mathsf{p}}}^{t_{N_{\mathsf{p}},j}},\quad t_{i,j}\in\mathbb{R},\quad\forall i=1,\dots,N_{\mathsf{p}},\quad\forall j=1,\dots,N_{\mathsf{x}},$ (A.3)
where the exponents $t_{i,j}$ in (A.3) are unknowns of scaling approaches.
Property A.4.
Property A.5.
Property A.6.
$[\tilde{p}_{j}]=\mathrm{u}_{1}^{m_{1,j}}\,\mathrm{u}_{2}^{m_{2,j}}\,\cdots\,\mathrm{u}_{N_{\mathsf{u}}}^{m_{N_{\mathsf{u}},j}}\neq 1,\quad m_{i,j}\in\mathbb{R},\quad\forall i=1,\dots,N_{\mathsf{u}},\quad\forall j=1,\dots,N_{\mathsf{p}},$ (A.4)
with $\mathrm{u}_{1}^{\beta_{1}}\,\mathrm{u}_{2}^{\beta_{2}}\,\cdots\,\mathrm{u}_{N_{\mathsf{u}}}^{\beta_{N_{\mathsf{u}}}}=1\Leftrightarrow\beta_{1}=\beta_{2}=\cdots=\beta_{N_{\mathsf{u}}}=0$, and all the exponents $m_{i,j}$ in (A.4) being explicitly known real numbers.
Aiming to minimize the number of involved quantities, one can also require the independence of the $N_{\mathsf{l}}$ coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13), or, equivalently, verify that the rank of the matrix made by the exponents in (A.1) and (A.2) is equal to $N_{\mathsf{l}}$, i.e.,
$\mathrm{rank}\left(\begin{pmatrix}c_{1,1}&\cdots&c_{1,N_{\mathsf{x}}}&d_{1,1}&\cdots&d_{1,N_{\mathsf{p}}}\\ c_{2,1}&\cdots&c_{2,N_{\mathsf{x}}}&d_{2,1}&\cdots&d_{2,N_{\mathsf{p}}}\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ c_{N_{\mathsf{l}},1}&\cdots&c_{N_{\mathsf{l}},N_{\mathsf{x}}}&d_{N_{\mathsf{l}},1}&\cdots&d_{N_{\mathsf{l}},N_{\mathsf{p}}}\end{pmatrix}\right)=N_{\mathsf{l}}.$ (A.5)
In Section 5, we show that Properties A.1, A.2, A.4, A.5 and A.6 hold when considering the projectile model presented in Section 2.2. For such a physical model, we also assume Property A.3, as explicitly indicated by (28). Though such an assumption can be motivated by its dimensional consistency, there exist physical models satisfying Properties A.1, A.2, A.4, A.5 and A.6, while violating Property A.3 (see, for example, the equation for energy density (2.49) in [18]).
Appendix B Theorems
Theorem B.1 enables the computation of numerical values $\theta_{1},\dots,\theta_{N_{\mathsf{x}}}$ of the characteristic constants $\tilde{\theta}_{1},\dots,\tilde{\theta}_{N_{\mathsf{x}}}$ that minimize the deviation $C_{U}$ from unity of the magnitudes of $\lambda_{i}$ (13), with $i\in U\subseteq\{1,\dots,N_{\mathsf{l}}\}$.
Theorem B.1.
Given Property A.1 of coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13), $U\subseteq\{1,\dots,N_{\mathsf{l}}\}$ and any fixed choice of $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$, the values of $(\theta_{1},\dots,\theta_{N_{\mathsf{x}}})\in(0,\infty)^{N_{\mathsf{x}}}$ attaining the global minimum of the function
$C_{U}(\theta_{1},\dots,\theta_{N_{\mathsf{x}}})\coloneqq\sum_{i\in U}\left(\log_{10}(\lambda_{i}(\tilde{\theta}_{1},\dots,\tilde{\theta}_{N_{\mathsf{x}}},\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}))\right)^{2},\quad C_{U}:(0,\infty)^{N_{\mathsf{x}}}\to[0,\infty),$ (B.1)
$\sum_{k=1}^{N_{\mathsf{x}}}\left(\sum_{i\in U}\,c_{i,k}\,c_{i,j}\right)\rho_{k}=-\sum_{i\in U}\,c_{i,j}\,\log_{10}(\kappa_{i}),\quad\forall j=1,\dots,N_{\mathsf{x}},$ (B.2)
where $\rho_{k}\coloneqq\log_{10}(\theta_{k})$ and the numerical value $\kappa_{i}\in(0,\infty)$ of each $\tilde{\kappa}_{i}=\tilde{\kappa}_{i}(\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}})$ is fixed.

### v2

@@ 5 A Case Study: the classical projectile model
First, we use the Theorems from Section 3.1.1 to formulate the projectile model with the minimal number of parameters. Assumptions˜4, 5 and 6 are satisfied by the $N_{\mathsf{p}}=4$ physical parameters $\tilde{p}_{1},\tilde{p}_{2},\tilde{p}_{3},\tilde{p}_{4}$ (6), with
$\tilde{p}_{1}]=[\tilde{g}]=\mathrm{m}\,\mathrm{s}^{-2},\quad[\tilde{p}_{2}]=[\tilde{R}]=\mathrm{m},\quad[\tilde{p}_{3}]=[\tilde{h}_{0}]=\mathrm{m},\quad[\tilde{p}_{4}]=[\tilde{v}_{0}]=\mathrm{m}\,\mathrm{s}^{-1},$ (42)
being $\mathrm{m}$ meters and $\mathrm{s}$ seconds. Then, Theorem 2 guarantees that there exist only $N_{\mathsf{0}}=N_{\mathsf{p}}-\mathrm{rank}(\textbf{M})=2$ independent dimensionless parameters, where M is the matrix
@@
and the $N_{\mathsf{0}}=2$ independent dimensionless parameters can be computed as in (21), i.e.,
@@
Assumptions˜1 and 2 are fulfilled by $N_{\mathsf{l}}=4$ coefficients $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ (5), with
@@
Then, by also bearing Assumption˜3 for $\tilde{\theta}_{1},\tilde{\theta}_{2}$ (3), i.e.,
@@
Theorem 3 guarantees that the coefficients $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ (5) can be written as power-law monomials of $\pi_{1},\pi_{2}$ (44) only, i.e.,
@@
Finally, Theorem 5 indicates under Assumptions˜1, 2, 3, 4, 5 and 6 that the vector
@@
of exponents in (48) must satisfy the linear system $\textbf{S}\,\vec{\alpha}=\vec{s}$ (26), which, for the projectile model, reads in row echelon form as
@@ 5.1 Optimally Scaled Coefficients
Theorem 7 allows solving the optimization problem (39) for any fixed choice of $\pi_{1},\pi_{2}$ (44), since one can set (50) as the constraint $\boldsymbol{\Sigma}\,\vec{\alpha}=\vec{\sigma}$ (30) to satisfy the assumption (32). In particular, the system (31) can be solved to find optimal values of $\vec{\alpha}$ (49) for plugging into (48). Then, the system (31) reads as
@@ 5.2 Generalized Optimal Scaling $(\mathtt{s}=0)$: Results
Table 1 states that such coefficients allow minimizing the variability $\mathop{\mathrm{max}}_{i}\lambda_{i}/\mathop{\mathrm{min}}_{i}\lambda_{i}$. In order to verify such a claim, we compare the values achieved by coefficients (54) with the obtained ones by the following approach suggested in the literature [Holmes2009_ND, Scaling_Langtangen]. Provided the values of physical parameters $\tilde{p}_{1},\tilde{p}_{2},\tilde{p}_{3},\tilde{p}_{4}$ (6), it is possible to impose $N_{\mathsf{x}}=2$ out of the $N_{\mathsf{l}}=4$ coefficients $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ (5) to be equal to $1$. Then, one aims to solve the resulting system of equations with respect to $\tilde{\theta}_{1}$ and $\tilde{\theta}_{2}$. Finally, the corresponding values of $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ (5) are obtained by plugging the computed $\tilde{\theta}_{1}$ and $\tilde{\theta}_{2}$ into (5).
@@
The Matlab solver $\mathrm{\it{ode45}}$ [ShampineReicheltODEMatlab] is employed to compute the numerical solution $\varphi=\varphi_{n}$ to the Ordinary Differential Equation (4) in the time interval $0,T/\theta_{1}]$. The subscript $n$ indicates that the numerical values of coefficients $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ and any computed quantity at each time step are rounded to $n$ decimal places. The relative error is estimated as
$e_{n}\coloneqq\frac{1}{T}\int_{0}^{T}\!\frac{\left|\theta_{2}\,\varphi_{n}(\tau/\theta_{1})-\varphi_{\mathrm{ref}}(\tau)\right|}{\left|\varphi_{\mathrm{ref}}(\tau)\right|}\,d\tau,$ (59)
where $\varphi_{\mathrm{ref}}$ is the numerical solution to the Ordinary Differential Equation (4) with $\lambda_{1}=g$, $\lambda_{2}=R^{-1}$, $\lambda_{3}=h_{0}$, $\lambda_{4}=v_{0}$, i.e., $\theta_{1}=\theta_{2}=1$, that has been taken as a reference in the time interval $[0,T]$. Such a solution $\varphi_{\mathrm{ref}}$ is obtained by using the Matlab solver $\mathrm{\it{ode45}}$ [ShampineReicheltODEMatlab] with the default 16 significant digits of precision for numerical computations. A relative error tolerance of $10^{-5}$ and an absolute error tolerance of $10^{-8}$ have been required for all the computations performed by $\mathrm{\it{ode45}}$ [ShampineReicheltODEMatlab], being $2$ orders of magnitude smaller than the default tolerances. The Simpson’s quadrature rule [DG.2024.simps] has been used to compute the integral in (59).
@@ 5.3 Generalized Optimal Scaling $(\mathtt{s}=1,2,3,4)$: Results
$\epsilon_{\mathtt{s}}(T)$ $\coloneqq\frac{\int_{0}^{T}\!\left|\varphi_{\mathtt{s}}(\tau)-\varphi_{\mathrm{ref}}(\tau)\right|^{2}\,d\tau}{\int_{0}^{T}\!\left|\varphi_{\mathrm{ref}}(\tau)\right|^{2}\,d\tau},\quad\mbox{with}\quad T\coloneqq\mathop{\mathrm{min}}\left\{T_{0},T_{\mathtt{m}}\right\},\quad T_{\mathtt{m}}\in(0,\infty),$
@@
where $\varphi_{\mathtt{s}}(\tau)$ is the approximating solution (60), while $\varphi_{\mathrm{ref}}(\tau)$ is the numerical solution to (4) provided by the Matlab solver $\mathrm{\it{ode45}}$ [ShampineReicheltODEMatlab], that has been taken as a reference since it is the solution one aims to approximate. If $\mathtt{s}=3,4$, the approximating solution (60) is also numerically computed by means of the Matlab solver $\mathrm{\it{ode45}}$ [ShampineReicheltODEMatlab]. A relative error tolerance of $10^{-6}$ and an absolute error tolerance of $10^{-9}$ have been required for all the computations performed by $\mathrm{\it{ode45}}$ [ShampineReicheltODEMatlab], i.e., $3$ orders of magnitude smaller than the default tolerances. The Simpson’s quadrature rule [DG.2024.simps] has been used to compute the integrals in (61). The time $T$ in (61) is chosen with the aim of stopping the numerical integration just before singularities are attained in (4) or (60). Such irregularities appear when the computed solutions of (4) or (60) are negative, i.e., equal to $1/\lambda_{2}<0$. In other words, the time interval $[0,T]$ in (61) allows investigating the dynamics of the projectile further than it reaches the ground level. The resulting error $\epsilon_{\mathtt{s}}(T)$ (61) is shown in Figure 4 as a function of $\pi_{1}$ and $\pi_{2}$ (44).
@@ 6 Conclusions & Discussion
We reviewed methodologies for scaling of dimensional models and presented a novel approach which we called Generalized Optimal Scaling, or GOS. GOS inherits all the features of its predecessors, i.e., Optimal Scaling (OS) and Optimal Scaling with Constraints (OSC), including the ability to quantify conditions for asymptotic models and, in addition, provides control of over-parametrization - the property beneficial for complex multiscale models. Both OS and OSC can be recovered from GOS with the special choices of the GOS’ parameter $\mathtt{s}$. Such a non-negative integer $\mathtt{s}$ is the index of the coefficient $\lambda_{\mathtt{s}}$ vanishing to 0 when an asymptotic model is considered. The GOS method was rigorously formulated and carefully tested on the classical projectile model.
By setting $\mathtt{s}=0$ in GOS (see Section 4), one makes use of the Optimal Scaling methodology [RUSCONI2019106944] presented in Section 3.2. For the projectile model, we demonstrated that OS allows reducing the variability of the involved coefficients by at least two more orders of magnitude compared to other scaling approaches traditionally applied to this model. Such a reduction is beneficial for diminishing round-off and numerical errors. Indeed, we showed that the coefficients provided by the Optimal Scaling approach allow reducing round-off errors by two more orders of magnitude compared to the traditional scaling techniques.
When GOS is applied with $\mathtt{s}\geq 1$ (see Section 4), one performs the Optimal Scaling with Constraints [RUSCONI2023127756] explained in Section 3.3. Such an approach provides a threshold on values of parameters that allow discarding terms in the considered equations, without significantly modifying the solutions. Thus, one can drop some terms and consider a simpler model, if the provided criterion is met. For the projectile model, OSC is able to identify values of parameters yielding to a relative approximation error of at most $10\%$ for all simulated intervals of time and all $\mathtt{s}\geq 1$.
We remark that GOS is applicable to models of arbitrary levels of complexity, which will be presented in upcoming studies. Moreover, a natural future step is an incorporation of GOS in an automated tool to streamline model reduction and scaling, optimizing parameter selection before further modeling steps. This would facilitate the use of GOS in complex, high-dimensional models, even for users with less mathematical expertise.
@@ Acknowledgements
This work was partially supported by a grant of the Ministry of Research, Innovation and Digitization, CNCS-UEFISCDI, project number PN-IV-P2-2.1-T-TE-2023-1704, within PNCDI IV.
Appendix A Proof of Theorem 1
See 1